Recent Log Entries
Crew Chief Shift Log, 2023-02-14 DAY
Start of shift: Beam to Halls A, B, C, and D.
Crew Chief Timesheet Summary: 14 Feb 2023 DAY
QuickPic - xB - Full Survey
Searching for 60 Hz noise in Injector
Took Control Access to search for 60 Hz noise in the Injector with EXTECH EMF probe but could not find any.
Used EXTECH Electromagnetic Field EMF ADAPTER Model 480824. The adapter was attached to a Voltmeter. It has a sensitivity of 1mGauss - 1mVdc and a range of 0.1mGuass to 199.9mGauss with 30 to 300Hz operational bandwidth.
The EMF noise was <0.2 mV everywhere in the vicinity of beamline - see for example Figure 1.
Initially, the Choppers were ON and the probe could is sensitive to this- see Figure 2.
Hall A Full Survey
Follow-up Re: NLinac heat
HYST AREA 6-2-4 Cycles Magnets for LEM
LEM South Linac
Gun HV Turned Off
Gun HV Rampdown Initiated
Hall C Full Survey
Hall D Full Survey
COMMENT ON: LAM3C Earth Leakage Fault Investigation
SBS Ramped Down
The OPS DRVHs for the following cavities have been updated: 1L16-8
The OPS DRVHs for the following cavities have been updated:
1L16-8
Username: benesch
Date/Time: 02-14-23_11:54:13
RF Status: Good
Tuner Status: Good
Comment: DRVH=5.350 Arc Rate Management five at 5.45, last four equally spaced 2h interval
Previous OPS DRVH: 5.45
Special Instructions:
The OPS DRVHs for the following cavities have been updated: 1L14-3
The OPS DRVHs for the following cavities have been updated:
1L14-3
Username: benesch
Date/Time: 02-14-23_11:51:51
RF Status: Good
Tuner Status: Good
Comment: DRVH=10.700 Arc Rate Management five at 10.8, equally spaced
Previous OPS DRVH: 10.8
Special Instructions:
The OPS DRVHs for the following cavities have been updated: 1L09-2
The OPS DRVHs for the following cavities have been updated:
1L09-2
Username: benesch
Date/Time: 02-14-23_11:50:36
RF Status: Good
Tuner Status: Good
Comment: DRVH=6.300 Arc Rate Management five at 6.35 irregular spacing
Previous OPS DRVH: 6.7
Special Instructions:
The OPS DRVHs for the following cavities have been updated: 1L20-6
The OPS DRVHs for the following cavities have been updated:
1L20-6
Username: benesch
Date/Time: 02-14-23_11:49:21
RF Status: Good
Tuner Status: Good
Comment: DRVH=6.600 Arc Rate Management seven at 6.8, equally spaced
Previous OPS DRVH: 6.8
Special Instructions:
LEM 9.16.1 testing
I completed full testing of the new LEM version. I performed various 'Apply' scenarios including slight adjustment to current profile, drop all cavities to minimal gradient, drop High Voltage on a zone, and pause a zone while others continue. I was unable to reproduce what Daniel had seen a while back when I first attempted to release this version.
I will install the new version soon and we should keep an eye out for any problems.